A circuit transforms 2.2×103 J of energy when it is operated for 3.0 min.Determine the amount of energy it will transform when it is operated for 1 h.
1
Expert's answer
2020-12-03T10:58:53-0500

Energy is directly proportional to time. "\frac{E}{E'} =\frac{t}{t'}"

"E'=22×103×\frac{60}{3}"

"E'=45320 J"